# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ III.3 Start-up shear
The success shown in Fig. 13 implies that the particle-level solid-liquid duality provides a potential scenario for understanding the flow of supercooled liquids. With above considerations, now we explore the response of supercooled liquids to the start-up shear. Our objective is to predict the stress evolution during the start-up process with the knowledge of the equilibrium state, such as the equilibrium spectrum of $\tau_{\mathrm{LC}}$. The form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is crucial here. As we discussed in Sec. II.4, for a specific sample, the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is given by Eq. 4 once a few material-dependent parameters are known. Note that for the study of amorphous solids, the deformation and plastic yielding in the start-up process are of central importance [86, 103]. However, studies focusing on the start-up process of supercooled liquids are much less. While some research [104] has identified the shear banding phenomenon in supercooled liquids under high shear rates, a more general understanding remains elusive.
@@
Borrowing some ideas from EPM, we construct a simple model for supercooled liquids under shear. We still assume that the system is separated into many cells. Nevertheless, there is a key difference between our model and EPM: In our model, the inter-cell elasticity-mediated interaction is not present, and cells are independent of each other. To understand this setting, we recall that in supercooled liquids, LERs are localized, and are separated by liquid-like regions [42, 83]. In this case, phonons can only propagate within LER, while are quickly damped in liquid-like regions [91]. Therefore, if we assume that a cell conceptually corresponds to an LER or a liquid-like region (see Fig. 14 (a) for an illustration), it will be unnecessary to consider the elasticity-mediated interaction between cells. Consequently, we assume that each cell evolves independently. Notice that in supercooled liquids, it is suggested that the relaxations of different dynamic regions are also coupled [109, 110]. To describe this effect, a coupling factor can be introduced [110]. At the current stage, we ignore this effect to keep the model as simple as possible.
As shown in Fig. 2, the spatial distribution of $\tau_{\mathrm{LC}}$ is smooth and tends to cluster in space. Thus, we can assign an average $\tau_{\mathrm{LC}}$ value to each cell. In the beginning of the shear, we sample the average $\tau_{\mathrm{LC}}$ value for each cell according to the equilibrium spectrum of $\tau_{\mathrm{LC}}$ ($H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$). Then, cells can be identified as fast or slow according to Eq. 2.
@@
To account for this thermal effect, after assigning each cell a $\tau_{\mathrm{LC}}$ sampled from $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$, we use the preceding distribution to sample $\tau_{\mathrm{LC}}^{\mathrm{th}}$ for each cell. With $\tau_{\mathrm{LC}}^{\mathrm{th}}$, the local dynamics under shear, represented by $\tau_{\mathrm{sh}}^{\mathrm{th}}$, can be found through Eq. 4.
@@
Now we discuss the evolution of cell. Slow cells behave as solid and can be viewed as LERs. As we discuss in Sec. III.1, the yielding and rearrangement of LER are collective due to the existence of the elasticity-mediated interaction. To embody this consideration in modeling, we recall our calculation in the preceding subsection: We use $\tau_{\mathrm{el}}$ as the duration time of the elastic stage of LER, rather than the average of $\tau_{\mathrm{sh}}$ of slow particles $\bar{\tau}_{\mathrm{sh,slow}}$. $\tau_{\mathrm{el}}$ is noticeably shorter than $\bar{\tau}_{\mathrm{sh,slow}}$. $\tau_{\mathrm{el}}$ corresponds to the waiting time for about $10\%$ of slow particles to relax. In contrast, $\bar{\tau}_{\mathrm{sh,slow}}$ represents the time at which nearly half of slow particles have relaxed. For such a long time, the slow region has long lost its elasticity. To capture this effect, we introduce the ratio $c_{\mathrm{el}}$:
@@
[…]
@@
In Fig. 15, we compare the start-up stress predicted by our model with the accurate results for both samples. In each column, we display the stress evolutions of slow particles, fast particles, and the whole system in sequence for a specific sample. Here, “slow” and “fast” denote the classification at the beginning of shear. It is seen that this simple model captures some essential features:
@@
There are some aspects that our model does not well align with the MD results, particularly at large shear rates. First, our model displays an undershoot following the overshoot. This undershoot arises from our assumption that $\tau_{\mathrm{sh}}^{\mathrm{th}}$ is solely determined from $\tau_{\mathrm{LC}}^{\mathrm{th}}$ through Eq. 4. In reality, there could be other uncertainties in this relation. By incorporating additional parameter, such as the yielding rate [105], this undershoot can be mitigated. More notably, the model overestimates the yielding strain for shear rates within the shear-thinning regime. This problem should be related to the fact that at high shear rates, the majority of the system behaves as solid-like regions. In this case, the behaviors of different cells are inevitably correlated by the elasticity-mediated interaction. As discussed above, the key difference between our model and EPM is that our model does not incorporate the correlation between different cells, while EPM does. Following this line of thought, we perform an EPM analysis [92, 105, 111, 116] on our samples. The details of the EPM we use will be given in a separated paper. In Fig. 15 (e), we plot the EPM predictions for sample A at $\dot{\gamma}=0.00025$ and $0.0011$. It is seen that for $\dot{\gamma}=0.0011$ (within the shear-thinning regime), the EPM gives a better prediction on the yielding strain than our model. However, at $\dot{\gamma}=0.00025$ (within the crossover regime), for which the solid-like behavior is not dominant, the performance of EPM is worse than our model. The reason could be that the inter-cell correlation mediated by elasticity, which is essential in EPM, is weak here due to the presence of massive liquid-like regions. In principle, EPM treats the mechanical behavior from the viewpoint of solids. On the contrary, the starting point of our model is the viewpoint of supercooled liquids—the heterogeneities in dynamics and response. Considering these results, a possible improvement of our model is to introduce the inter-cell correlation between adjacent solid-like cells. This work is in progress.
@@ IV STRUCTURAL BASIS
The above analysis shows that the intricate flow behaviors of supercooled liquids can be understood by focusing on $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$, which are connected by the relation $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ given by Eq. 4. There are several key observations related to $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$: First, $\tau_{\mathrm{LC}}$ exhibits DH features, in the sense that $\tau_{\mathrm{LC}}$ exhibits a broad distribution and significant clustering in space. Second, the functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ of a liquid is universal once the final shear rate is given. To be specific, it converts a $\tau_{\mathrm{LC}}$ value to the corresponding $\tau_{\mathrm{sh}}$, regardless of the configuration from which $\tau_{\mathrm{LC}}$ is extracted. The degeneracy of $\tau_{\mathrm{LC}}$ implies that there exists an underlying configurational parameter (or a set of parameters) that can be applied to both equilibrium and shear states to determine the $\tau_{\mathrm{LC}}$ value through the same functional form. Third, as the final shear rate increases, $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ bends at smaller $\tau_{\mathrm{LC}}$ with more remarkable flattening. This change directly leads to the increase of the solid-like component in flow, which critically enhances shear thinning. Here, a question arises: How can we understand these dynamical behaviors within a unified framework?
As we mentioned in Sec. I, in the last two decades, substantial efforts have been made to explore the role of structure in determining dynamics of supercooled liquids and glasses. Inspired by this thinking, in this section, we aim to seek the structural basis of the dynamical behaviors mentioned in the preceding paragraph. The key is to construct a parameter or a set of parameters using the structural properties of equilibrium state and shear states with various $\dot{\gamma}$. By employing this structural parameter, we reveal a connection between microscopic structure and dynamics for flowing supercooled liquids. This connection, together with the established relation between $\tau_{\mathrm{LC}}$ and rheology, creates a pathway from microscopic structure to macroscopic flow behavior.
The discussion in this section is based on the condition of steady shear. The conclusions can also be applied to the start-up shear.
@@ IV.1 Structural parameters for $\tau_{\mathrm{LC}}$
We first establish the connection between $\tau_{\mathrm{LC}}$ and configuration. In literature of supercooled liquids and glasses, numerous structural parameters have been proposed to predict dynamics, ranging from simple, physically informed metrics to more complex ones. Some examples include free volume [36, 117], local favored structures [37, 45], soft modes [25], structural entropy [35], etc. In recent years, approaches based on statistical learning have gained prominence [28, 32, 33, 34]. Especially, some sophisticated models, such as graph neural network [33, 118] and multilayer perceptron [119], have been employed and achieved unprecedentedly high prediction accuracies compared with the single, simpler parameters, albeit at the cost of being order-agnostic. Notably, linear regression method offers a balance between high prediction accuracy and interpretability. Another advantage of linear regress method is that it does not need a very large data set. In previous studies [34, 120], Filion et al. have connected the local structure of reference particle to its dynamic propensity using linear regression, which achieves an accuracy comparable to that of more complex models. In this subsection, linear regression will be used to connect $\tau_{\mathrm{LC}}$ and local configuration.
@@
Figure 16: Correlation between $\tau_{\mathrm{LC}}$ and the local structural feature defined by Eq. 9. (a) and (b) give the results of sample A and sample B, respectively. The correlation is characterized by the Pearson correlation coefficient $p$. The horizontal axis represents the element in the feature vector. Each element is characterized by $l$, $m$ and $r$ (only big particles are considered here). $l$ ranges from 00 to $6$, dividing the horizontal coordinates into $7$ regimes separated by vertical dashed lines. For each $l$, the values of $m$ are arranged from $-l$ to $l$. Different $m$ features with the same $l$ are distinguished by different colors. For each $(l,m)$ combination, there are $48$ values of $r$. In total, there are $2352$ elements are shown. Two prominent peaks are highlighted by black arrows, representing the features constructed using $g_{0}^{0}$ and $g_{2}^{-2}$.
@@
For a given combination of $(l,m)$, the feature vector $g_{l}^{m}$ comprises $96$ elements, involving $48$ $r$ values and $2$ species. In principle, one may include numerous $(l,m)$ combinations to formulate the total feature vector for a particle. However, too many $(l,m)$ combinations would yield a feature vector that is cumbersome and computationally intensive. To make our feature representation as simple as possible, it is necessary to judiciously select a subset of $(l,m)$ combinations to serve as the major features. For this purpose, we examine the importance of every $(l,m)$ combination in affecting dynamics up to $l=6$ by computing the Pearson correlation coefficient $p$ between each element in the feature vector with respect to the given $(l,m)$ and the target property $\tau_{\mathrm{LC}}$. For the sake of simplicity, only big particles are considered here. Thus, $48$ elements, covering all $r$ values, are examined for a $(l,m)$ combination. Considering that the number of $(l,m)$ combinations up to $l=6$ is $49$, we examine $48\times 49=2352$ elements in total for each sample. Figure 16 shows the results. The horizontal axis represents the feature index ranging from $1$ to $2352$. These indexes are divided into $7$ regimes according to the value of $l$. For each $l$, the values of $m$ are arranged from $-l$ to $l$; and for each $(l,m)$ combination, the values of $r$ are arranged from $0.3$ to $5.0$ in Fig. 16. Two prominent peaks in Fig. 16 are seen, respectively corresponding to $(l=0,m=0)$ and $(l=2,m=-2)$. The $(l=0,m=0)$ features reflect the packing efficiency around the reference particle, and it has been reported that local packing capacity is highly correlated with dynamics [26, 29]. The $(l=2,m=-2)$ features have a symmetry consistent with the shear geometry [42, 114, 115], and thus reflect the local shear strain [42, 72, 78, 121]. As discussed in the previous sections, different $\tau_{\mathrm{LC}}$ values correspond to different deformation phases in solid-like regions. A larger local strain indicates that it is closer to the local yielding point, leading to greater instability. Thus, the high correlation of $(l=2,m=-2)$ features is expected. In passing, the importance of the local packing and configurational distortion in the dynamics and rheology of supercooled liquids has been highlighted in recent studies [122, 123]. It is also seen that features corresponding to $l=6$ show relatively high correlation. This is consistent with the studies reporting the existence of icosahedral pattern in supercooled liquids, which is strongly correlated with dynamics [124, 125].
@@
According to the observation shown in Fig. 16, we use the features with $(l=0,m=0)$ and $(l=2,m=-2)$ to construct the feature vector. For particle $i$, the feature vector is given by:
@@
Incorporating both species and $48$ $r$ values, the elements of $X_{i}^{(0)}$ amount to $192$. The superscript “$(0)$” denotes that this is the zeroth generation of a series of feature vectors.
@@
$X_{i}^{(n+1)}=\frac{1}{Z}{\textstyle\sum_{j,r_{ij}<r_{c}}\exp(-\frac{r_{ij}}{r_{c}})X_{j}^{(n)}},$ (11)
where $r_{c}$ is a parameter characterizing the range of coarse-graining and is set to $4$, and $Z$ is written as:
$Z={\textstyle\sum_{j,r_{ij}<r_{c}}\exp(-\frac{r_{ij}}{r_{c}})}.$ (12)
Another advantage of the coarse-graining is that it averages out irrelevant thermal noise. With Eqs. 10 and 11, we construct the final feature vector for particle $i$ by combining the feature vectors with generations 00, $1$ and $2$:
@@
Figure 18: Cumulative Pearson correlation coefficients of the top $30$ most important elements in the feature vector defined by Eq. 13. (a) and (b) show the results of sample A and sample B, respectively. The dashed line in each panel denotes the Pearson correlation coefficient obtained with the entire feature vector.
We randomly split our data set with a ratio of $8:2$, with $80\%$ of the data used for training and the remaining $20\%$ used for testing. Figure 17 compares the true values of $\ln\tau_{\mathrm{LC}}$ and the values predicted by the regression model for the test set, where each point represents a particle. Here, results under steady shear with various $\dot{\gamma}$, including equilibrium, Newtonian, crossover, and shear-thinning regimes, are plotted together. These results suggest a positive correlation between the predicted and true values of $\ln\tau_{\mathrm{LC}}$, quantified by the Pearson correlation coefficient $p$ of $0.87$ for sample A, and $0.82$ for sample B. We also calculate the Pearson correlation coefficient for each $\dot{\gamma}$. The results, given in the insets of Fig. 17, are around $0.7$. It is seen that the correlation for single $\dot{\gamma}$ remains robust. The good correlations shown in Fig. 17 suggest that the structural parameters used here are applicable to both equilibrium and a wide range of $\dot{\gamma}$.
An important observation from Fig. 17 is that the data points corresponding to different $\dot{\gamma}$ approximately align along the same truth line. This consistency suggests the existence of a common structural basis for $\tau_{\mathrm{LC}}$ that is invariant to the applied shear. As a result, we unify the $\tau_{\mathrm{LC}}$ found from equilibrium and shear states by leveraging these structural parameters.
One may argue that the $576$ elements contained in one feature vector given by Eq. 13 are still too many, which calls for further interpretation and simplification. We sort these elements according to their contributions to the cumulative Pearson correlation coefficient. Figure 18 gives the cumulative Pearson correlation coefficients of the top $30$ most important elements in sorting sequence. It is seen that the first several elements contribute the majority of the total correlation, while the remaining ones’ contribution is marginal. The detailed sequences of important elements for both samples are given in Appendix C.

### v2

@@ III.3 Start-up shear
The success shown in figure 13 implies that the particle-level solid-liquid duality provides a potential scenario for understanding the flow of supercooled liquids. With above considerations, now we explore the response of supercooled liquids to the start-up shear. Our objective is to predict the stress evolution during the start-up process with the knowledge of the equilibrium state, such as the equilibrium distribution of $\tau_{\mathrm{LC}}$. The form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is crucial here. As we discussed in section II.4, for a specific sample, the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ is given by equation 4 once a few material-dependent parameters are known. Note that for the study of amorphous solids, the deformation and plastic yielding in the start-up process are of central importance [86, 103]. However, studies focusing on the start-up process of supercooled liquids are much less. While some research [104] has identified the shear banding phenomenon in supercooled liquids under high shear rates, a more general understanding remains elusive.
@@
Borrowing some ideas from EPM, we construct a simple model for supercooled liquids under shear. We still assume that the system is separated into many cells. Nevertheless, there is a key difference between our model and EPM: In our model, the inter-cell elasticity-mediated interaction is not present, and cells are independent of each other. To understand this setting, we recall that in supercooled liquids, LERs are localized, and are separated by liquid-like regions [44, 83]. In this case, phonons can only propagate within LER, while are quickly damped in liquid-like regions [91]. Therefore, if we assume that a cell conceptually corresponds to an LER or a liquid-like region (see figure 14 (a) for an illustration), it will be unnecessary to consider the elasticity-mediated interaction between cells. Consequently, we assume that each cell evolves independently. Notice that in supercooled liquids, it is suggested that the relaxations of different dynamic regions are also coupled [49, 109]. To describe this effect, a coupling factor can be introduced [109]. At the current stage, we ignore this effect to keep the model as simple as possible.
As shown in figure 2, the spatial distribution of $\tau_{\mathrm{LC}}$ is smooth and tends to cluster in space. Thus, we can assign an average $\tau_{\mathrm{LC}}$ value to each cell. In the beginning of the shear, we sample the average $\tau_{\mathrm{LC}}$ value for each cell according to the equilibrium distribution of $\tau_{\mathrm{LC}}$ value ($H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$). Then, cells can be identified as fast or slow according to equation 2.
@@
To account for this thermal effect, after assigning each cell a $\tau_{\mathrm{LC}}$ sampled from $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$, we use the preceding distribution to sample $\tau_{\mathrm{LC}}^{\mathrm{th}}$ for each cell. With $\tau_{\mathrm{LC}}^{\mathrm{th}}$, the local dynamics under shear, represented by $\tau_{\mathrm{sh}}^{\mathrm{th}}$, can be found through equation 4.
@@
Now we discuss the evolution of cell. Slow cells behave as solid and can be viewed as LERs. As we discuss in section III.1, the yielding and rearrangement of LER are collective due to the existence of the elasticity-mediated interaction. To embody this consideration in modeling, we recall our calculation in the preceding subsection: We use $\tau_{\mathrm{el}}$ as the duration time of the elastic stage of LER, rather than the average of $\tau_{\mathrm{sh}}$ of slow particles $\bar{\tau}_{\mathrm{sh,slow}}$. $\tau_{\mathrm{el}}$ is noticeably shorter than $\bar{\tau}_{\mathrm{sh,slow}}$. $\tau_{\mathrm{el}}$ corresponds to the waiting time for about $10\%$ of slow particles to relax. In contrast, $\bar{\tau}_{\mathrm{sh,slow}}$ represents the time at which nearly half of slow particles have relaxed. For such a long time, the slow region has long lost its elasticity. To capture this effect, we introduce the ratio $c_{\mathrm{el}}$:
@@
In fast cells, the elasticity spanning an extended spatial region is not present. Cage jump is highly independent, in the sense that the phonon propagation induced by such a flow event is very short-lived and localized [91]. Therefore, we can put our attention on the cage. Under shear, the cage of a fast particle slightly deforms [111, 112] and rearranges quickly after a jump time $\tau_{\mathrm{sh}}$. The massive and successive cage jumps in a fast region result in a stress evolution that is relatively stationary, without an overall stress accumulation process. To model the cycle of fast cells, however, we still assume a deformation–yielding process, as illustrated in figure 14 (c). This form should not be viewed as the evolution of a solid-like cluster. On the contrary, it represents the averaged cage distortion–cage jump process in a fast cell. Notice that, here we still use $G_{\mathrm{eff}}$ as the slope of stress, as shown in figure 14 (c). In principle, the modulus here should be smaller than that of solid-like regions. However, considering that modulus is not very sensitive to relaxation time [80], we adopt $G_{\mathrm{eff}}$ for fast cells as an acceptable approximation. By assuming a similar apparent form of cycle for both fast and slow cells, we simplify the calculation of our model.
@@
In figure 15, we compare the start-up stress predicted by our model with the accurate results for both samples. In each column, we display the stress evolutions of slow particles, fast particles, and the whole system in sequence for a specific sample. Here, “slow” and “fast” denote the classification at the beginning of shear. It is seen that this simple model captures some essential features:
@@
There are some aspects that our model does not well align with the MD results, particularly at large shear rates. First, our model displays an undershoot following the overshoot. This undershoot arises from our assumption that $\tau_{\mathrm{sh}}^{\mathrm{th}}$ is solely determined from $\tau_{\mathrm{LC}}^{\mathrm{th}}$ through equation 4. In reality, there could be other uncertainties in this relation. By incorporating additional parameter, such as the yielding rate [105], this undershoot can be mitigated. More notably, the model overestimates the yielding strain for shear rates within the shear-thinning regime. This problem should be related to the fact that at high shear rates, the majority of the system behaves as solid-like regions. In this case, the behaviors of different cells are inevitably correlated by the elasticity-mediated interaction. As discussed above, the key difference between our model and EPM is that our model does not incorporate the correlation between different cells, while EPM does. Following this line of thought, we perform an EPM analysis [92, 105, 110, 113] on our samples. The details of the EPM we use will be given in a separated paper. In figure 15 (e), we plot the EPM predictions for sample A at $\dot{\gamma}=0.00025$ and $0.0011$. It is seen that for $\dot{\gamma}=0.0011$ (within the shear-thinning regime), the EPM gives a better prediction on the yielding strain than our model. However, at $\dot{\gamma}=0.00025$ (within the crossover regime), for which the solid-like behavior is not dominant, the performance of EPM is worse than our model. The reason could be that the inter-cell correlation mediated by elasticity, which is essential in EPM, is weak here due to the presence of massive liquid-like regions. In principle, EPM treats the mechanical behavior from the viewpoint of solids. On the contrary, the starting point of our model is the viewpoint of supercooled liquids—the heterogeneities in dynamics and response. Considering these results, a possible improvement of our model is to introduce the inter-cell correlation between adjacent solid-like cells. This work is in progress.
@@ IV STRUCTURAL BASIS
The above analysis shows that the intricate flow behaviors of supercooled liquids can be understood by focusing on $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$, which are connected by the relation $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ given by equation 4. There are several key observations related to $\tau_{\mathrm{sh}}$ and $\tau_{\mathrm{LC}}$: First, $\tau_{\mathrm{LC}}$ exhibits DH features, in the sense that $\tau_{\mathrm{LC}}$ exhibits a broad distribution and significant clustering in space. Second, the functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ of a liquid is universal once the final shear rate is given. To be specific, it converts a $\tau_{\mathrm{LC}}$ value to the corresponding $\tau_{\mathrm{sh}}$, regardless of the configuration from which $\tau_{\mathrm{LC}}$ is extracted. The degeneracy of $\tau_{\mathrm{LC}}$ implies that there exists an underlying configurational parameter (or a set of parameters) that can be applied to both equilibrium and shear states to determine the $\tau_{\mathrm{LC}}$ value through the same functional form. Third, as the final shear rate increases, $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ bends at smaller $\tau_{\mathrm{LC}}$ with more remarkable flattening. This change directly leads to the increase of the solid-like component in flow, which critically enhances shear thinning. Here, a question arises: How can we understand these dynamical behaviors within a unified framework?
As we mentioned in section I, in the last two decades, substantial efforts have been made to explore the role of structure in determining dynamics of supercooled liquids and glasses. Inspired by this thinking, in this section, we aim to seek the structural basis of the dynamical behaviors mentioned in the preceding paragraph. The key is to construct a parameter or a set of parameters using the structural properties of equilibrium state and shear states with various $\dot{\gamma}$. By employing this structural parameter, we reveal a connection between microscopic structure and dynamics for flowing supercooled liquids. This connection, together with the established relation between $\tau_{\mathrm{LC}}$ and rheology, creates a pathway from microscopic structure to macroscopic flow behavior.
@@ IV.1 Structural parameters for $\tau_{\mathrm{LC}}$
The discussion in this section will be based on the condition of steady shear. First, we will establish the connection between $\tau_{\mathrm{LC}}$ and configuration. In literature of supercooled liquids and glasses, numerous structural parameters have been proposed to predict dynamics, ranging from simple, physically informed metrics to more complex ones. Some examples include free volume [38, 114], local favored structures [39, 47], soft modes [27], structural entropy [37], etc. In recent years, approaches based on statistical learning have gained prominence [30, 34, 35, 36]. Especially, some sophisticated models, such as graph neural network [35, 115] and multilayer perceptron [116], have been employed and achieved unprecedentedly high prediction accuracies compared with the single, simpler parameters, albeit at the cost of being order-agnostic. Notably, linear regression method offers a balance between high prediction accuracy and interpretability. Another advantage of linear regress method is that it does not need a very large data set. In previous studies [36, 117], Filion et al. have connected the local structure of reference particle to its dynamic propensity using linear regression, which achieves an accuracy comparable to that of more complex models. In this subsection, linear regression will be used to connect $\tau_{\mathrm{LC}}$ and local configuration.
@@
Figure 16: Correlation between $\tau_{\mathrm{LC}}$ and the local structural feature defined by equation 9. (a) and (b) give the results of sample A and sample B, respectively. The correlation is characterized by the Pearson correlation coefficient pp. The horizontal axis represents the element in the feature vector. Each element is characterized by $l$, $m$ and $r$ (only big particles are considered here). $l$ ranges from 0 to $6$, dividing the horizontal coordinates into $7$ regimes separated by vertical dashed lines. For each $l$, the values of $m$ are arranged from l-l to $l$. Different $m$ features with the same $l$ are distinguished by different colors. For each $(l,m)$ combination, there are $48$ values of $r$. In total, there are $2352$ elements are shown. Two prominent peaks are highlighted by black arrows, representing the features constructed using $g_{0}^{0}$ and $g_{2}^{-2}$.
@@
For a given combination of $(l,m)$, the feature vector $g_{l}^{m}$ comprises $96$ elements, involving $48$ $r$ values and $2$ species. In principle, one may include numerous $(l,m)$ combinations to formulate the total feature vector for a particle. However, too many $(l,m)$ combinations would yield a feature vector that is cumbersome and computationally intensive. To make our feature representation as simple as possible, it is necessary to judiciously select a subset of $(l,m)$ combinations to serve as the major features. For this purpose, we examine the importance of every $(l,m)$ combination in affecting dynamics up to $l=6$ by computing the Pearson correlation coefficient $p$ between each element in the feature vector with respect to the given $(l,m)$ and the target property $\tau_{\mathrm{LC}}$. For the sake of simplicity, only big particles are considered here. Thus, $48$ elements, covering all $r$ values, are examined for a $(l,m)$ combination. Considering that the number of $(l,m)$ combinations up to $l=6$ is $49$, we examine $48\times 49=2352$ elements in total for each sample. Figure 16 shows the results. The horizontal axis represents the feature index ranging from $1$ to $2352$. These indexes are divided into $7$ regimes according to the value of $l$. For each $l$, the values of $m$ are arranged from l-l to $l$; and for each $(l,m)$ combination, the values of $r$ are arranged from $0.3$ to $5.0$ in figure 16. Two prominent peaks in figure 16 are seen, respectively corresponding to $(l=0,m=0)$ and (l=2,m=2)(l=2,m=-2). The $(l=0,m=0)$ features reflect the packing efficiency around the reference particle, and it has been reported that local packing capacity is highly correlated with dynamics [28, 31]. The (l=2,m=2)(l=2,m=-2) features have a symmetry consistent with the shear geometry [44, 111, 112], and thus reflect the local shear strain [44, 72, 78, 118]. As discussed in the previous sections, different $\tau_{\mathrm{LC}}$ values correspond to different deformation phases in solid-like regions. A larger local strain indicates that it is closer to the local yielding point, leading to greater instability. Thus, the high correlation of (l=2,m=2)(l=2,m=-2) features is expected. In passing, the importance of the local packing and configurational distortion in the dynamics and rheology of supercooled liquids has been highlighted in recent studies [119, 120]. It is also seen that features corresponding to $l=6$ show relatively high correlation. This is consistent with the studies reporting the existence of icosahedral pattern in supercooled liquids, which is strongly correlated with dynamics [121, 122].
@@
According to the observation shown in figure 16, we use the features with $(l=0,m=0)$ and (l=2,m=2)(l=2,m=-2) to construct the feature vector. For particle $i$, the feature vector is given by:
@@
Incorporating both species and $48$ $r$ values, the elements of $X_{i}^{(0)}$ amount to $192$. The superscript “(0)(0)” denotes that this is the zeroth generation of a series of feature vectors.
@@
$X_{i}^{(n+1)}=\frac{1}{Z}{\textstyle\sum_{j,r_{ij}<r_{\mathrm{c}}}\exp(-\frac{r_{ij}}{r_{\mathrm{c}}})X_{j}^{(n)}},$ (11)
where $r_{\mathrm{c}}$ is a parameter characterizing the range of coarse-graining and is set to $4$, and $Z$ is written as:
$Z={\textstyle\sum_{j,r_{ij}<r_{\mathrm{c}}}\exp(-\frac{r_{ij}}{r_{\mathrm{c}}})}.$ (12)
Another advantage of the coarse-graining is that it averages out irrelevant thermal noise. With equations 10 and 11, we construct the final feature vector for particle $i$ by combining the feature vectors with generations 0, 11 and $2$:
@@
Figure 18: Cumulative Pearson correlation coefficients of the top $30$ most important elements in the feature vector defined by equation 13. (a) and (b) show the results of sample A and sample B, respectively. The dashed line in each panel denotes the Pearson correlation coefficient obtained with the entire feature vector.
We randomly split our data set with a ratio of $8:2$, with $80\%$ of the data used for training and the remaining $20\%$ used for testing. Figure 17 compares the true values of $\ln\tau_{\mathrm{LC}}$ and the values predicted by the regression model for the test set, where each point represents a particle. Here, results under steady shear with various $\dot{\gamma}$, including equilibrium, Newtonian, crossover, and shear-thinning regimes, are plotted together. These results suggest a positive correlation between the predicted and true values of $\ln\tau_{\mathrm{LC}}$, quantified by the Pearson correlation coefficient $p$ of $0.87$ for sample A, and $0.82$ for sample B. We also calculate the Pearson correlation coefficient for each $\dot{\gamma}$. The results, given in the insets of figure 17, are around $0.7$. It is seen that the correlation for single $\dot{\gamma}$ remains robust. The good correlations shown in figure 17 suggest that the structural parameters used here are applicable to both equilibrium and a wide range of $\dot{\gamma}$.
An important observation from figure 17 is that the data points corresponding to different $\dot{\gamma}$ approximately align along the same truth line. This consistency suggests the existence of a common structural basis for $\tau_{\mathrm{LC}}$ that is invariant to the applied shear. As a result, we unify the $\tau_{\mathrm{LC}}$ found from equilibrium and shear states by leveraging these structural parameters.
One may argue that the $576$ elements contained in one feature vector given by equation 13 are still too many, which calls for further interpretation and simplification. We sort these elements according to their contributions to the cumulative Pearson correlation coefficient. Figure 18 gives the cumulative Pearson correlation coefficients of the top $30$ most important elements in sorting sequence. It is seen that the first several elements contribute the majority of the total correlation, while the remaining ones’ contribution is marginal. The detailed sequences of important elements for both samples are given in Appendix C.
Figure 19: Spatial distributions of the hardness $H$ at various $\dot{\gamma}$ for both samples. (a) – (d) show the configuration slices selected at different $\dot{\gamma}$ for sample A. (e) – (h) show the configuration slices selected at different $\dot{\gamma}$ for sample B. In all panels, the color of each particle denotes its value of $H$ with a common color bar. The boundaries between liquid-like (fast) and solid-like (slow) groups are denoted by black dashed lines. A nice agreement between the boundaries and the particles with $H$ close to 0 is seen in all panels.